OpenStudy (anonymous):

Last problem I need help with! Please! Using the given data, find the values of sine, cosine, and tangent of 2 A and the quadrant in which 2 A terminates: angle A in Quadrant II, cos A = -7/25 cos 2 A = a. 527/625 b. -336/625 c. 336/625 d. -527/625

OpenStudy (ajprincess):

\(\cos2A=2\cos A\sin A\) Find the value of sinA using \(\sin A=\sqrt{1-\cos^2A}\). Can u do nw?

OpenStudy (anonymous):

So 2(-0.28)(0.276)

OpenStudy (anonymous):

:/ It's not matching up

OpenStudy (ajprincess):

Find them in the fraction form. Vat do u get for sinA?

OpenStudy (anonymous):

69/250

OpenStudy (anonymous):

cos=Adjacent/Hypotenuse and sin=Opposite/Hypotenuse and tan=Opposite/ajacent so here you have given just adjacent and hypotenuse and you can fine opposite by Pythagourus Theorum.

OpenStudy (ajprincess):

\(\sin A=\sqrt{1-\cos^2A}\) \(\sin A=\sqrt{1-(\large\frac{-7}{25})^2}\) \(\sin A=\sqrt{1-\large\frac{49}{625}}\) \(\sin A=\sqrt{\large\frac{625-49}{625}}\) \(\sin A=\sqrt{\large\frac{576}{625}}\) \(\sin A=\large\frac{24}{25}\) Nw can u find cos2A?

OpenStudy (anonymous):

its hard method friend. find sin by sin=perp/hyp we need perp so from cos=base/hyp here base=-7 and hyp=25 so now use Pythagoras theorem to find perp.

OpenStudy (anonymous):

Yes! -336/625. Thank you so much for step by step explanation! @ajprincess Thank you to @muhammad9t5 as well. I appreciate the time and effort my friends

OpenStudy (ajprincess):

U r welcome.:)

OpenStudy (anonymous):

hyp^2=base^+perp^2 25^2=-7^2+perp^2 625=49+perp^2 625-49=perp^2 576=perp^2 24=perp

OpenStudy (anonymous):

thanks! you are always welcome friend.

OpenStudy (ajprincess):

I agree vth u @muhammad9t5. Ur method is easy.

OpenStudy (anonymous):

yup. @ajprincess please fan me.

OpenStudy (ajprincess):

I fanned u. i fan all those who fan me.:)